# Multi-Microphone Speech Emotion Recognition using the Hierarchical Token-semantic Audio Transformer Architecture

### v2

@@ 3.2 Architecture
We will now elaborate on the HTS-AT architecture. The standard transformer architecture requires extensive computational resources due to the unmodified input token sequence length across all layers. This includes maintaining a large global self-attention matrix and calculating outputs and gradients at each step. The HTS-AT architecture is introduced to address these challenges. The HTS-AT is designed for supporting multiple audio tasks, e.g., classification, sound event detection, and source localization. It introduces two critical architectural optimizations: a hierarchical structure and a windowed attention mechanism. The input audio mel-spectrogram is split into localized patch tokens using a convolutional Patch-Embed CNN layer of kernel dimensions $P\times P$, in which the patches are ordered by time segment and frequency bin. Then, tokens propagate through a series of Swin Transformer encoder groups. At the terminus of each group, a Patch Merging layer reshapes the token sequence into its original 2D mel-spectrogram. This layer merges neighboring patches and then embeds them back into a latent space of reduced length. Consequently, the memory requirements decrease exponentially with depth. Within each Swin Transformer block, attention is restricted to non-overlapping $M\times M$ squares, partitioning the token sequence. Calculating self-attention within each window subset substantially reduces computational complexity relative to full global attention while capturing localized relationships. As patch size increases downstream, windows encapsulate larger temporal and frequency contexts. Finally, the HTS-AT incorporates a token-semantic CNN layer after the last transformer block. This layer refines the output by grouping tokens, thus capturing information about their time frames and frequency bins. Consequently, this enhances classification by exploiting token relationships.
@@ 3.3 Multi-Channel Methods
Fig. 1: Scheme of patch tokens mel-spectrogram segments.
Fig. 2: Scheme of average mel-spectrogram.
Patch-Embed Scheme: Inspired by the study in [29], we embraced a Patch-Embed scheme where each channel undergoes a shared embedding layer followed by summation. This maintains flexibility in the microphone numbers used during fine-tuning and testing without altering the core HTS-AT architecture. On the left side of Fig. 1, each of the $M$ channels is analyzed to generate mel-spectrograms, which are then concatenated along the channel depth. SpecAugment [30] is applied collectively across all $M$ mel-spectrograms. Subsequently, each mel-spectrogram is reshaped to the dimensions of a $256\times 256$ image and passes through the shared Patch-Embed layer. Next, a summation operation is performed across all $M$-encoded channels, making this input suitable for the pre-trained HTS-AT. The summation consolidates inter-channel information, enabling more robust representations suitable for reverberant conditions. The summed embeddings are then fed to the HTS-AT feedforward block for the actual emotion classification.
@@ 4.1 Datasets
We used the gpuRIR Python package222github.com/DavidDiazGuerra/gpuRIR to simulate reverberant multi-channel microphone signals (four channels) with reverberation time in the range of $T_{60}=200-800$ ms with different room sizes and randomized microphone locations. The room dimensions were uniformly distributed between 3 and 8 meters, with an aspect ratio ranging from 1 to 1.6. The room’s height was fixed at 2.9 meters. The sound source was placed at an elevation of 1.75 meters, while the microphones were positioned at 1.6 meters. Both the sound source and microphones were kept at least 0.5 meters away from the surrounding walls. By doing so, we were also able to enlarge the datasets. For RAVDESS, 6863 reverberant speech samples were generated for training, and 852 samples were used for validation. We follow the same procedure for IEMOCAP with 7356 train samples and 2107 validation samples and for CREMA-D with 5945 train samples and 1487 validation samples. Our objective encompasses evaluating our model in real-world reverberant environments. To achieve this, we conducted tests utilizing the ACE RIR database [26], which comprises seven distinct rooms characterized by varying dimensions and exhibiting diverse ranges of $T_{60}$. We used the subset captured by a mobile phone equipped with three microphones, which are ubiquitous and often used for speech interaction, making them a practical choice for real-world SER applications. The model was fine-tuned with the synthesized RIRs and evaluated with the test sets of the various datasets convolved with the ACE RIRs.
@@
Table 2: Results on the test sets of the RAVDESS, IEMOCAP, and CREMA-D datasets convolved with RIRs of three microphones from the ACE database. The ‘HTS-AT’ columns are fine-tuned on reverberant single-channel audio. The ‘Avg mel’ columns depict results where mel-spectrograms were averaged across four channels during fine-tuning and tested on three channels. The ‘Sum PE’ columns are the Patch-Embed Scheme approach fine-tuned and tested on three channels.
@@ 4.3 Results
We now turn to the evaluation of the multi-channel schemes. In all experiments, we tested all three datasets with the utterances convolved with the three RIRs of the ACE database. We only report the results for the remote source case to emphasize the reverberation effects. The average mel-spectrogram scheme was fine-tuned using four microphones. The Patch-Embed scheme was fine-tuned using three microphones. While we used different microphone counts (3 and 4) for fine-tuning, this was done to demonstrate the flexibility of our approach in handling varying numbers of input channels. However, we acknowledge that using the same number of microphones for all models would provide a more direct comparison. In both cases, the RIRs were generated using the simulator, as explained above. All experimental procedures followed identical training-validation splits, maintaining consistent model configurations, sizes, and hyperparameters for each dataset.
Table LABEL:ACE_results provides a comparative assessment of three model variants, all fine-tuned with reverberant speech: 1) ‘HTS-AT’ - a single-channel scheme, 2) ‘Avg mel’ - mel-spectrograms averaged across four channels, and 3) ‘Sum PE’ - the Patch-Embed scheme. In the table, we report the accuracy and the associated confidence interval for all seven rooms and three databases. Examining these results indicates the consistent, albeit modest, improvements achieved by the ‘Sum PE’ approach. Embedding each channel separately before summing, the model captures subtle inter-channel differences that might be lost in averaging. This allows the model to emphasize the most informative aspects of each channel for emotion recognition, leading to improved performance. Importantly, these benefits come without a considerable increase in computational complexity compared to single-channel models.

### v3

@@ 3.2 Architecture
We will now elaborate on the HTS-AT architecture. The standard transformer architecture requires extensive computational resources due to the unmodified input token sequence length across all layers. This includes maintaining a large global self-attention matrix and calculating outputs and gradients at each step. The HTS-AT architecture is introduced to address these challenges. The HTS-AT is designed for supporting multiple audio tasks, e.g., classification, sound event detection, and source localization. It introduces two critical architectural optimizations: a hierarchical structure and a windowed attention mechanism. The input audio mel-spectrogram is split into localized patch tokens using a convolutional Patch-Embed CNN layer of kernel dimensions $P\times P$, in which the patches are ordered by time segment and frequency bin. Then, tokens propagate through a series of Swin Transformer encoder groups. At the terminus of each group, a Patch Merging layer reshapes the token sequence into its original 2D mel-spectrogram. This layer merges neighboring patches and then embeds them back into a latent space of reduced length. Consequently, the memory requirements decrease exponentially with depth. Within each Swin Transformer block, window attention is restricted to non-overlapping $W\times W$ squares, partitioning the token sequence. Calculating self-attention within each window subset substantially reduces computational complexity relative to full global attention while capturing localized relationships. As patch size increases downstream, windows encapsulate larger temporal and frequency contexts. Finally, the HTS-AT incorporates a token-semantic CNN layer after the last transformer block. This layer refines the output by grouping tokens, thus capturing information about their time frames and frequency bins. Consequently, this enhances classification by exploiting token relationships.
@@ 3.3 Multi-Channel Methods
Fig. 1: Scheme of Patch-Embed Summation.
Fig. 2: Scheme of Average Mel-Spectrograms.
Patch-Embed Summation: Inspired by the study in [29], we embraced a Patch-Embed scheme where each channel undergoes a shared embedding layer followed by summation. This maintains flexibility in the microphone numbers used during fine-tuning and testing without altering the core HTS-AT architecture. On the left side of Fig. 1, each of the $M$ channels is analyzed to generate mel-spectrograms, which are then concatenated along the channel depth. SpecAugment [30] is applied collectively across all $M$ mel-spectrograms. Subsequently, each mel-spectrogram is reshaped to the dimensions of a $256\times 256$ image and passes through the shared Patch-Embed layer. Next, a summation operation is performed across all $M$-encoded channels, making this input suitable for the pre-trained HTS-AT. The summation consolidates inter-channel information, enabling more robust representations suitable for reverberant conditions. The summed embeddings are then fed to the HTS-AT feedforward block for the actual emotion classification.
@@ 4.1 Datasets
We used the ’gpuRIR’ Python package222github.com/DavidDiazGuerra/gpuRIR to simulate reverberant multi-channel microphone signals (setting the number of microphones to $M=4$) with reverberation time in the range of $T_{60}=200-800$ ms with different room sizes and randomized microphone locations. The room dimensions were uniformly distributed between 3 and 8 meters, with an aspect ratio ranging from 1 to 1.6. The room’s height was fixed at 2.9 meters. The sound source was placed at a constant height of 1.75 meters, while the microphones were positioned at a fixed height of 1.6 meters. The $x$ and $y$ coordinates were randomly determined within the room for both the sound source and microphones, while keeping them at least 0.5 meters away from the room walls. By doing so, we were also able to enlarge the datasets. For RAVDESS, 6863 reverberant speech samples were generated for training, and 852 samples were used for validation. We follow the same procedure for IEMOCAP with 7356 train samples and 2107 validation samples and for CREMA-D with 5945 train samples and 1487 validation samples. Our objective encompasses evaluating our model in real-world reverberant environments. To achieve this, we conducted tests utilizing the ACE RIR database [26], which comprises seven distinct rooms characterized by varying dimensions and exhibiting diverse ranges of $T_{60}$. We used the subset captured by a mobile phone equipped with three microphones, which are frequently used for speech interaction, making them a practical choice for real-world SER applications. The model was fine-tuned with the synthesized RIRs and evaluated with the test sets of the various datasets convolved with the ACE RIRs.
@@
Table 2: Results on the test sets of the RAVDESS, IEMOCAP, and CREMA-D datasets convolved with RIRs of three microphones from the ACE database. The ‘HTS-AT’ columns are fine-tuned on reverberant single-channel audio. The ‘Avg mel’ columns depict results where mel-spectrograms were averaged across four channels during fine-tuning and tested on three channels. The ‘Sum PE’ columns are the Patch-Embed Summation approach fine-tuned and tested on three channels.
@@ 4.3 Results
We now turn to the evaluation of the multi-channel schemes. In all experiments, we tested all three datasets with the utterances convolved with the three RIRs of the ACE database. We only report the results for the remote source case to emphasize the reverberation effects. The average mel-spectrogram scheme was fine-tuned using four microphones. The Patch-Embed Summation scheme was fine-tuned using three microphones. While we used different microphone counts ($M=3$ and $M=4$) for fine-tuning, this was done to demonstrate the flexibility of our approach in handling varying numbers of input channels. However, we acknowledge that using the same number of microphones for all models would provide a more direct comparison. In both cases, the RIRs were generated using the simulator, as explained above. All experimental procedures followed identical training-validation splits, maintaining consistent model configurations, sizes, and hyperparameters for each dataset.
Table LABEL:ACE_results provides a comparative assessment of three model variants, all fine-tuned with reverberant speech: 1) ‘HTS-AT’ - a single-channel scheme, 2) ‘Avg mel’ - mel-spectrograms averaged across four channels, and 3) ‘Sum PE’ - Patch-Embed Summation across three channels. In the table, we report the accuracy and the associated confidence interval for all seven rooms and three databases. Examining these results indicates the consistent, albeit modest, improvements achieved by the ‘Sum PE’ approach. Embedding each channel separately before summing, the model captures subtle inter-channel differences that might be lost in averaging. This allows the model to emphasize the most informative aspects of each channel for emotion recognition, leading to improved performance. Importantly, these benefits come without a considerable increase in computational complexity compared to single-channel models.